OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation x - 4y = 4. slope = 4; y intercept at (0, -1) slope = ; y intercept at (0, 1) slope = ; y intercept at (0, -1) slope = -4; y intercept at (0, 1

OpenStudy (stamp):

y = mx + b Right now you have x - 4y = 4 Get your equation into the form y = mx + b Once you have arrived, identify your slope = m and your y-intercept = b